30 ml of 1 M NaOH reacts with 10 ml 1M H2SO4 - how many moles of Na2SO4 are produced & what is the limiting reactant ?
Answer #1

2NaOH + H​​​​​​2​​​​​SO 4   Na​​2SO4 + 2H2O

Moles of NaOH = molarity×volume/1000

= 1×30/1000 = 0.03 moles

Moles of H2SO4 = 1×10/1000 = 0.01 moles

According to the reaction,

2 moles of NaOH require 1 mole H2SO4

0.03 moles of NaOH require 0.03/2 = 0.015 moles of H2SO4

Since we have less than this, hence H2SO4 is a limiting reagent.

According to the reaction,

1 mole H2SO4 give 1 mole Na2SO4

0.01 mole of H2SO4 give 0.01 moles of Na2SO4

Please give feedback!